Determine the interest rate earned on a $1950 deposit when $2250 is paid back in one year
Answer #1

Answer:

PV = FV/(1+r)^n
PV - Present value
FV - Future value
r - Interest rate
n - no. of periods
1950 = 2250/(1+r)^1
r = (2250/1950)^1 - 1 = 0.15846 = 15.846%
Rate of interest =15.85%